Grade 11General Physics

If the error in measurements of velocity of a particle is 10% and mass is known exactly. The percentage error in K.E is?

Ajay Mittur
10 Years agoGrade 11

1 Answer

Aryan Gupta
10 Years ago
If you mean that the mass measured was 100% precise (i.e. no error or 0% error),the percentage error in the K.E. is 20%.Solution:error in mass = 0%error in velocity = 10%K.E. = m*v*vtherefore % error in K.E. = 0+10+10= 20%.regardsand please approve my answer!